OpenStudy (anonymous):

2(x + 2) + 8 = 6 Solve this equation and show all work.

OpenStudy (kobeni-chan):

First, you want to use the distributive property on 2(x+2). Multiply x by 2, and +2 by 2. Then, add 8, and let me know what you have.

OpenStudy (anonymous):

2x+4+8=6

OpenStudy (kobeni-chan):

That's right! Now you can combine the 4 and 8 to make it 2x + 12 = 6. The next step is to get the 2x part by itself, on one side of the equal sign. To do this, we can subtract 12 from both sides. Do 2x + 12 - 12 = 6 - 12

OpenStudy (anonymous):

2x=6 so x= 3

OpenStudy (kobeni-chan):

Right! :)

OpenStudy (kobeni-chan):

oh wait!! 6 - 12 is -6, so it will actually be -2

OpenStudy (anonymous):

-3?

OpenStudy (kobeni-chan):

-3 yes. Sorry about that confusion lol

OpenStudy (anonymous):

its okay thanks for the help

OpenStudy (kobeni-chan):

No problem!